# Q-Mamba: On First Exploration of Vision Mamba for Image Quality Assessment

### v1

@@ Abstract
In this work, we take the first exploration of the recently popular foundation model, i.e., State Space Model/Mamba, in image quality assessment, aiming at observing and excavating the perception potential in vision Mamba. A series of works on Mamba has shown its significant potential in various fields, e.g., segmentation and classification. However, the perception capability of Mamba has been under-explored. Consequently, we propose Q-Mamba by revisiting and adapting the Mamba model for three crucial IQA tasks, i.e., task-specific, universal, and transferable IQA, which reveals that the Mamba model has obvious advantages compared with existing foundational models, e.g., Swin Transformer, ViT, and CNNs, in terms of perception and computational cost for IQA. To increase the transferability of Q-Mamba, we propose the StylePrompt tuning paradigm, where the basic lightweight mean and variance prompts are injected to assist the task-adaptive transfer learning of pre-trained Q-Mamba for different downstream IQA tasks. Compared with existing prompt tuning strategies, our proposed StylePrompt enables better perception transfer capability with less computational cost. Extensive experiments on multiple synthetic, authentic IQA datasets, and cross IQA datasets have demonstrated the effectiveness of our proposed Q-Mamba.
footnotetext: ${}^{\textrm{\Letter}}$ Corresponding authors.
@@ 1 Introduction
Image Quality Assessment (IQA) aims to measure the subjective quality of images aligned with human perception, which has been applied in various visual fields, including visual acquisition, transmission, AIGC [25, 61, 63, 71], and UGC creation [58, 31, 42], etc. Establishing a great IQA metric is necessary to provide the right optimization direction tailored for image processing techniques, i.e., compression [30, 64, 8, 35], enhancement [27, 28, 24, 7], and ensure the perceptual quality of images. Early works on IQA have been achieved by leveraging the natural scene statistics in a hand-crafted manner [44, 46, 60, 52]. With the advancements of deep neural networks (DNNs), learning-based IQA metrics [29, 38, 34, 76, 77, 3] have demonstrated significant potential for low-level perception, which can be roughly categorized into two types based on the pre-traiend backbones: CNN-based and Transformer-based methods.
@@
Although the impressive progress, learning-based IQA are susceptible to inherent limitations of existing pre-trained backbones: (i) the CNNs are skilled at learning local translation-invariant features from images while lacking enough long-range dependency modeling capability, hindering the global quality perception. (ii) The emergence of Vision Transformers presents a great solution to model long-range dependency effectively by leveraging attention mechanisms. However, the quadratic complexity of self-attention operations poses unaffordable computational costs, especially for large-scale image quality assessment. Recently, an innovative foundation model, the State Space Model, particularly its implementation, i.e., Mamba [12] has shown considerable potential in various fields for balancing the computational costs and performances, e.g., segmentation[65, 43] and classification[54, 80]. This raises one interesting question, i.e., “whether Mamba can surpass existing backbones on low-level visual perception", which is still under-explored.
@@
To answer this question, in this work, we initiate a new exploration of the Mamba model within the field of image quality assessment (IQA), and introduce the Q-Mamba, a newly developed IQA framework designed to address three key facets of IQA: task-specific, universal, and transferable image quality assessment. Notably, since the lack of enough IQA dataset, learning-based IQA metrics entail the pre-trained backbone for perception knowledge extraction. Meanwhile, VMamba [39], as the most representative framework, has achieved excellent performance on high-level tasks by employing horizontal and vertical scanning strategies. However, merely excavating the global perception knowledge is not optimal for IQA, since most artifacts affecting the image quality are related to local textures. Consequently, inspired by LocalMamba[18], we have adopted a local window scanning method, which significantly enhances Mamba’s ability to perceive local distortions, thereby demonstrating superior performance. We detail the architecture of Q-Mamba and examine different scanning methods to clarify how the model interacts with and processes image data. Our extensive analysis confirms that Q-Mamba surpasses traditional foundational models, showing superior perceptual accuracy and greater computational efficiency.
@@
Moreover, we have explored the perception transferability of Mamba across different datasets, i.e., different contents, and degradations. We can find that the Mamba-based IQA metric still suffers from severe performance drop when they encounter large domain shifts between synthetic distortions[53, 33], authentic distortions[10, 17], and Artificial Intelligence-Generated Content (AIGC)[25, 61] distortions, etc. To further amplify the transferability of Q-Mamba across a range of IQA applications, we introduce a simple but effective tuning strategy called StylePrompt. This is based on the finding that the domain shifts in IQA tend to correlate with their feature statistics/style [41], such as mean and variance. Concretely, our StylePrompt aims to adaptively adjust the mean and variance of pre-trained Q-Mamba towards the target IQA tasks by setting a group of light-weight learnable $1\times 1\times C$ parameters. Extensive experiments have shown that our StylePrompt greatly facilitates the task-adaptive learning of Q-Mamba, enabling efficient knowledge transfer across diverse IQA tasks with fewer parameter costs.
@@
  • We embark on a novel exploration of the Mamba model within image quality assessment (IQA), and propose the Q-Mamba, a powerful IQA metric for three critical tasks of IQA: task-specific, universal, and transferable image quality assessment. This exploration has demonstrated the superior potential of Mamba for subjective perception, advancing the development of IQA.
  • To improve the perception transferability of the Q-Mamba, we introduce a simple but effective tuning strategy, i.e., StylePrompt. This strategy enable the efficient knowledge transfer of pre-trained Q-Mamba for downstream IQA tasks, while only tuning fewer learnable parameters to adjust the statistics of perception features.
  • Extensive experiments have shown that our Q-Mamba has consistently achieved state-of-the-art results on various prominent IQA datasets compared with existing IQA methods, thereby validating the efficacy of the Mamba model in quality assessment. Moreover, our StylePrompt achieves nearly equivalent performance to full model tuning while utilizing only $4\%$ of the whole parameters, demonstrating its effectiveness in perception knowledge transfer scenarios.
@@ 2.1 Blind Image Quality Assessment (BIQA)
Early BIQA methods relied heavily on manually designed features for quality score regression [44, 46, 60, 52]. However, these handcrafted features were insufficient for addressing the complexity of BIQA tasks. With the advent of deep learning, network architectures capable of powerful feature extraction significantly improved quality assessment tasks, with Convolutional Neural Networks (CNNs) and Vision Transformers being the most prevalent.
CNN-based BIQA. CNNs have demonstrated robust feature extraction capabilities, leading to their widespread adoption in BIQA tasks. Early works like CNNIQA [20] used convolutional models for feature learning and quality regression, substantially outperforming handcrafted features. DBCNN [75] introduced a dual-stream network to address synthetic and authentic distortions separately, integrating these insights for better quality prediction. NIMA [57] and PQR [72] leveraged pre-trained models on ImageNet for quality score prediction, enhancing accuracy through well-established neural architectures. MetaIQA [79] employed meta-learning to adapt to unknown distortions by learning shared priors for various distortion types, while HyperIQA [56] used a hypernetwork to adaptively establish perceptual rules, improving generalization. Despite these advancements, CNNs’ local bias limits their ability to fully exploit both global and local information in BIQA tasks.
Transformer-based BIQA. Transformers offer superior global modeling capabilities compared to CNNs. TReS [11] addressed CNNs’ local bias by capturing local structural information with CNNs and then using Transformers for sequential feature extraction. MUSIQ [21] designed a multi-scale image Transformer architecture capable of handling images with varying sizes and aspect ratios. DEIQT [50] leveraged a Transformer-based BIQA architecture with attention mechanisms to align with human perception, enhancing model performance and reducing prediction uncertainty. However, the quadratic complexity of Transformers presents a challenge, highlighting the need for architectures with linear complexity capable of global modeling for BIQA tasks.
@@ 2.2 State Space Models (SSMs)
State-space models, known for their linear complexity in capturing long-range dependencies, have been integrated into deep learning architectures. The Structured State-Space model (S4) [14] was a pioneer in deep state-space modeling for remote dependency modeling. Subsequent advancements [13, 55, 9] further propelled the development of state-space models. Mamba [12], by integrating selection mechanisms and hardware-aware algorithms, has shown effective long-range modeling capabilities with linear complexity growth.
Initially focused on NLP tasks, Mamba has rapidly expanded into other domains. Vim [80] introduced a bidirectional SSM block for visual representation learning, achieving performance comparable to ViT [5]. VMamba [39] introduced a cross-scan module to traverse spatial domains and convert non-causal visual images into ordered block sequences, maintaining linear complexity while retaining global receptive fields. LocalMamba [18] employed a window-based scanning approach to integrate local inductive biases, enhancing the visual Mamba model. These advancements validate Mamba’s efficacy in visual tasks, leading to its application in image classification [39, 80, 47, 67, 1], video understanding [62, 4, 26], image restoration [15, 54, 78], point cloud analysis [32, 74, 37], and biomedical image segmentation [43, 65, 51, 36]. This paper explores Mamba’s potential in visual quality perception tasks, providing a simple yet effective baseline for future quality perception research.
@@ 3 Method
3.1 Prelimineries
State Space Model (SSM).
Traditional State Space Models (SSMs) utilize first-order differential equations to establish a linear time-invariant system, translating one-dimensional input sequences $x(t)\in\mathbb{R}$ into outputs through hidden states $h(t)$ and $y(t)$, described by:
$\begin{gathered}h^{\prime}(t)=Ah(t)+Bx(t)\\ y(t)=Ch(t)\end{gathered}$ (1)
For compatibility with deep learning’s sequential input requirements, these models undergo a transition to a discrete framework within the S4 model, using discretization parameters $\Delta$ to modify $A$ and $B$ into $\overline{A}$ and $\overline{B}$:
$\begin{gathered}h_{t}=\overline{A}h_{t-1}+\overline{B}x_{t}\\ y_{t}=Ch_{t}\end{gathered}$ (2)
$\begin{gathered}\overline{A}=\exp(\Delta A)\\ \overline{B}=(\Delta A)^{-1}(\exp(\Delta A)-I)\cdot\Delta B\end{gathered}$ (3)
Mamba[12], evolving from these foundations, introduces a selection algorithm for dynamic adaptation of $A$, $B$, and $C$ based on input, enhancing model response to complex sequences and contextual variations. However, this necessitates abandoning traditional convolutional methods for kernel fusion and parallel scanning to maintain computational efficiency, effectively scaling with sequence length and providing advanced capabilities in sequence generation for language models.
Figure 1: Overview of the Q-Mamba architecture and description of the StylePrompt.
@@ 3.2.1 Overall Framework
The overall architecture of our Q-Mamba model is illustrated in Figure1. To successfully incorporate the Mamba model into visual perception tasks, we implemented the Q-Mamba architecture following the construction method of the VMamba’s visual Mamba blocks. This architecture consists of multiple layers of VMamba blocks, with the SS2D block serving as the core module, which facilitates information transmission through a residual network. The residual network incorporates convolutional layers for feature extraction and activation layers for computing gating signals. We employed multiple network stages, each comprising a downsampling layer and a VMamba Block, which construct multi-level representations at varying resolutions through downsampling, thus extracting richer perceptual features. To explore the impact of model size on quality perception tasks, we experimented with three different scales of the model: QMamba-Tiny, QMamba-Small, and QMamba-Base.
@@ 3.2.2 Perception with Local Scanning
Originally, Mamba demonstrated exceptional proficiency in NLP tasks, characterized by inherently causal sequential inputs. However, when adapted to visual tasks, which lack corresponding causal frameworks, the model encountered substantial difficulties. It struggled to effectively discern the complex interdependencies among image pixels, thereby impeding its learning processes.
VMamba[39] employs a novel bidirectional horizontal and vertical scanning mechanism, illustrated in Figure 2(a), to convert two-dimensional data into one-dimensional inputs for model ingestion. Although this method facilitates global modeling of pixel sequences, it unfortunately causes locally adjacent tokens to be spaced significantly apart, thus failing to capture essential local dependencies. This limitation becomes particularly acute in quality assessment tasks, where accurate distortion detection is dependent on the granularity of local detail.
Drawing inspiration from LocalMamba[18], we have implemented a window-based scanning strategy, as depicted in Figure 2(b), to enhance local perceptual accuracy. This involves conducting horizontal scans within each window followed by comprehensive window-level horizontal scans, with a parallel approach adopted in the vertical dimension. This method markedly enhances VMamba’s capability to detect local quality details. Despite the model’s input sequence undergoing downsampling and aggregation, this scanning approach retains an ability for global perception, ensuring a robust analytical balance that effectively captures both detailed local nuances and broader contextual information.
Figure 2: Scanning Methodology Illustration. (a) VMamba’s[39] method flattens 2D data into 1D, impairing connectivity by distancing adjacent tokens. (b) Our local scanning method[18] scans within and across windows, placing semantically similar and distortion-related tokens closer, as shown in the blue boxes.

### v2

# QMamba: On First Exploration of Vision Mamba for Image Quality Assessment

@@ Abstract
In this work, we take the first exploration of the recently popular foundation model, i.e., State Space Model/Mamba, in image quality assessment (IQA), aiming at observing and excavating the perception potential in vision Mamba. A series of works on Mamba has shown its significant potential in various fields, e.g., segmentation and classification. However, the perception capability of Mamba remains under-explored. Consequently, we propose QMamba by revisiting and adapting the Mamba model for three crucial IQA tasks, i.e., task-specific, universal, and transferable IQA, which reveals its clear advantages over existing foundational models, e.g., Swin Transformer, ViT, and CNNs, in terms of perception and computational cost. To improve the transferability of QMamba, we propose the StylePrompt tuning paradigm, where lightweight mean and variance prompts are injected to assist task-adaptive transfer learning of pre-trained QMamba for different downstream IQA tasks. Compared with existing prompt tuning strategies, our StylePrompt enables better perceptual transfer with lower computational cost. Extensive experiments on multiple synthetic, authentic IQA datasets, and cross IQA datasets demonstrate the effectiveness of our proposed QMamba. The code will be available at: https://github.com/bingo-G/QMamba.git
Machine Learning, ICML
{guanfb, yuzihao, luyt31415}@mail.ustc.edu.cn, {xin.li, chenzhibo}@ustc.edu.cn
Figure 1: Scanning Methodology Illustration. (a) VMamba’s scanning method(Liu et al., 2024c) flattens 2D data into 1D, impairing connectivity by distancing adjacent tokens. (b) Local scanning method scans within and across windows, placing semantically similar and distortion-related tokens closer, as shown in the blue boxes.
@@ 1 Introduction
Image Quality Assessment (IQA) aims to measure the subjective quality of images aligned with human perception, which has been applied in various visual fields, including visual acquisition, transmission, AIGC (Li et al., 2023a; Wang et al., 2023a), and UGC creation (Tu et al., 2021; Lu et al., 2024), etc. Establishing a great IQA metric is necessary to provide the right optimization direction tailored for image processing techniques, i.e., compression (Li et al., 2021; Wu et al., 2021; Yu et al., 2024b), enhancement (Li et al., 2023b; Fei et al., 2023), and ensure the perceptual quality of images. Early works on IQA have been achieved by leveraging the natural scene statistics in a hand-crafted manner (Mittal et al., 2012a, b). With the advancements of deep neural networks (DNNs), learning-based IQA metrics (Li et al., 2023c; Liu et al., 2022) have demonstrated significant potential for low-level perception, which can be roughly categorized into two types based on the pre-trained backbones: CNN-based and Transformer-based methods.
@@
Although the impressive progress, learning-based IQA is susceptible to inherent limitations of existing pre-trained backbones: (i) the CNNs are skilled at learning local translation-invariant features from images while lacking enough long-range dependency modeling capability, hindering the global quality perception. (ii) The emergence of Vision Transformers presents a great solution to model long-range dependency effectively by leveraging attention mechanisms. However, the quadratic complexity of self-attention operations poses unaffordable computational costs, especially for large-scale image quality assessment. Recently, an innovative foundation model, the State Space Model, particularly its implementation, i.e., Mamba (Gu & Dao, 2023) has shown considerable potential in various fields for balancing the computational costs and performances, e.g., segmentation(Xing et al., 2024; Ma et al., 2024) and classification(Liu et al., 2024c; Zhu et al., 2024). This raises one interesting question, i.e., “whether Mamba can surpass existing backbones on low-level visual perception”, which is still under-explored.
@@
To answer this question, in this work, we initiate a new exploration of the Mamba model within the field of IQA and introduce the QMamba, a newly developed IQA framework designed to address three key facets of IQA: task-specific, universal, and transferable image quality assessment. Notably, since the lack of enough IQA dataset, learning-based IQA metrics entail the pre-trained backbone for perception knowledge extraction. Meanwhile, VMamba(Liu et al., 2024c), as the most representative framework, has achieved excellent performance on high-level tasks by employing horizontal and vertical scanning strategies. However, merely excavating the global perception knowledge is not optimal for IQA, since most artifacts affecting the image quality are related to local textures. Consequently, inspired by LocalMamba(Huang et al., 2024), we have adopted a local window scanning method, which significantly enhances ability of Mamba to perceive local distortions, thereby demonstrating superior performance. We detail the architecture of QMamba and examine different scanning methods to clarify how the model interacts with and processes image data. Our extensive analysis confirms that QMamba surpasses traditional foundational models, showing superior perceptual accuracy and greater computational efficiency.
@@
Moreover, we have explored the perceptual transferability of Mamba across different datasets, i.e., different contents, and degradations. We can find that the Mamba-based IQA metric still suffers from severe performance drop when they encounter large domain shifts between synthetic distortions(Sheikh et al., 2006; Lin et al., 2019), authentic distortions(Ghadiyaram & Bovik, 2015; Hosu et al., 2020), and Artificial Intelligence-Generated Content (AIGC)(Li et al., 2023a; Wang et al., 2023a) distortions, etc. To further amplify the transferability of QMamba across a range of IQA applications, we introduce a simple but effective tuning strategy called StylePrompt. This is based on the finding that the domain shifts in IQA tend to correlate with their feature statistics/style (Lu et al., 2022), such as mean and variance. Concretely, our StylePrompt aims to adaptively adjust the mean and variance of pre-trained QMamba towards the target IQA tasks by setting a group of light-weight learnable $1\times 1\times C$ parameters. Extensive experiments have shown that our StylePrompt greatly facilitates the task-adaptive learning of QMamba, enabling efficient knowledge transfer across diverse IQA tasks with fewer parameter costs.
@@
  • We embark on a novel exploration of the Mamba model within image quality assessment, and propose the QMamba, a powerful IQA metric for three critical tasks of IQA: task-specific, universal, and transferable image quality assessment. This exploration has demonstrated the superior potential of Mamba for subjective perception, advancing the development of IQA.
  • To improve the perception transferability of the QMamba, we introduce a simple but effective tuning strategy, i.e., StylePrompt. This strategy enables the efficient knowledge transfer of pre-trained QMamba for downstream IQA tasks, while only tuning fewer learnable parameters to adjust the statistics of perception features.
  • Extensive experiments have shown that our QMamba has consistently achieved state-of-the-art results on various prominent IQA datasets compared with existing IQA methods, thereby validating the efficacy of the Mamba model in quality assessment. Moreover, our StylePrompt achieves nearly equivalent performance to full model tuning while utilizing only $4\%$ of the whole parameters, demonstrating its effectiveness in perception knowledge transfer scenarios.
@@ 2.1 Blind Image Quality Assessment (BIQA)
Early BIQA methods relied heavily on manually designed features for quality score regression (Mittal et al., 2012a, b; Venkatanath et al., 2015; Saad et al., 2012; Min et al., 2018). However, these handcrafted features were insufficient for addressing the complexity of BIQA tasks. With the advent of deep learning, network architectures capable of powerful feature extraction significantly improved quality assessment tasks, with Convolutional Neural Networks (CNNs) and Vision Transformers being the most prevalent.
CNN-based BIQA. CNNs have demonstrated robust feature extraction capabilities, leading to their widespread adoption in BIQA tasks. Early works like CNNIQA (Kang et al., 2014) used convolutional models for feature learning and quality regression, substantially outperforming handcrafted features. DBCNN (Zhang et al., 2020) introduced a dual-stream network to address synthetic and authentic distortions separately, integrating these insights for better quality prediction. NIMA (Talebi & Milanfar, 2018) and PQR (Zeng et al., 2017) leveraged pre-trained models on ImageNet for quality score prediction, enhancing accuracy through well-established neural architectures. MetaIQA (Zhu et al., 2020) employed meta-learning to adapt to unknown distortions by learning shared priors for various distortion types, while HyperIQA (Su et al., 2020) used a hypernetwork to adaptively establish perceptual rules, improving generalization. Despite these advancements(Saha et al., 2023; Zhao et al., 2023; Chen et al., 2024a), CNNs’ local bias limits their ability to fully exploit both global and local information in BIQA tasks.
Figure 2: Architectural Overview of QMamba Framework and the Detailed of StylePrompt Tuning Mechanism.
Transformer-based BIQA. Transformers offer superior global modeling capabilities compared to CNNs. TReS (Golestaneh et al., 2022) addressed CNNs’ local bias by capturing local structural information with CNNs and then using Transformers for sequential feature extraction. MUSIQ (Ke et al., 2021) designed a multi-scale image Transformer architecture capable of handling images with varying sizes and aspect ratios. DEIQT (Qin et al., 2023) leveraged a Transformer-based BIQA architecture with attention mechanisms to align with human perception, enhancing model performance and reducing prediction uncertainty. However, the quadratic complexity of Transformers presents a challenge, highlighting the need for architectures with linear complexity capable of global modeling for BIQA tasks. In addition to the above methods, many recent works (Zhang et al., 2023; Shin et al., 2024; Yu et al., 2024a) have also adopted Transformer-based designs and achieved competitive performance, yet they still face limitations due to high computational complexity.
@@ 2.2 State Space Models (SSMs)
State space models, known for their linear complexity in capturing long-range dependencies, have been integrated into deep learning architectures. The Structured State Space model (S4) (Gu et al., 2021) was a pioneer in deep state space modeling for remote dependency modeling. Subsequent advancements (Gu et al., 2020; Smith et al., 2022; Fu et al., 2022) further propelled the development of state space models. Mamba (Gu & Dao, 2023), by integrating selection mechanisms and hardware-aware algorithms, has shown effective long-range modeling capabilities with linear complexity growth.
Initially focused on NLP tasks, Mamba has rapidly expanded into other domains. Vim (Zhu et al., 2024) introduced a bidirectional SSM block for visual representation learning, achieving performance comparable to ViT (Dosovitskiy et al., 2020). VMamba (Liu et al., 2024c) introduced a cross-scan module to traverse spatial domains and convert non-causal visual images into ordered block sequences, maintaining linear complexity while retaining global receptive fields. LocalMamba (Huang et al., 2024) employed a window-based scanning approach to integrate local inductive biases, enhancing the visual Mamba model. These advancements validate the efficacy of Mamba in visual tasks, leading to its application in image classification (Liu et al., 2024c; Zhu et al., 2024; Patro & Agneeswaran, 2024), video understanding (Wang et al., 2023b; Chen et al., 2024b; Li et al., 2024), image restoration (Guo et al., 2024; Shi et al., 2024; Zhen et al., 2024), point cloud analysis (Liang et al., 2024; Zhang et al., 2024; Liu et al., 2024b), and biomedical image segmentation (Ma et al., 2024; Xing et al., 2024; Liu et al., 2024a). These studies have demonstrated the effectiveness of state space models in visual tasks, providing a solid foundation for further exploration of their potential in visual perception.
@@ 3.1.1 Overall Framework
The overall architecture of our proposed QMamba model is depicted in Figure 2. To address the unique challenges of visual perception tasks, we developed a novel architecture that rethinks the design principles of state space models for visual data. The network incorporates a hierarchical residual structure, where convolutional layers enable effective feature extraction while specialized activation layers compute adaptive gating signals. Our architecture organizes the processing into multiple network stages, each combining a strategic downsampling layer with our enhanced Mamba-based processing block. This design enables the construction of multi-level representations at varying resolutions, facilitating the extraction of richer perceptual features through progressive abstraction.
To systematically investigate the relationship between model capacity and quality perception performance, we developed three distinct variants of our architecture: QMamba-Tiny, QMamba-Small, and QMamba-Base. Each variant maintains the core architectural principles while scaling in complexity, allowing us to explore the trade-offs between computational efficiency and perceptual accuracy across different application scenarios.
@@ 3.1.2 Perception with Local Scanning
While the original Mamba model excels in natural language processing tasks with inherently causal inputs, it faces challenges when applied to visual tasks due to the absence of spatial causality. In particular, it struggles to capture complex spatial dependencies among image pixels, which hinders its ability to model local distortions effectively.
To address this, VMamba (Liu et al., 2024c) introduces a bidirectional horizontal and vertical scanning strategy to convert 2D images into 1D sequences suitable for sequential modeling. Although this enables global pixel-level modeling, it disrupts the continuity of locally adjacent tokens, weakening the model’s ability to perceive fine-grained distortions—an essential factor for image quality assessment.
Inspired by LocalMamba (Huang et al., 2024), we adopt a window-based scanning approach that performs horizontal scans within local windows, followed by window-level scans, and applies the same strategy vertically. This method enhances local distortion perception while retaining global awareness through hierarchical composition. Despite downsampling and aggregation in the input, this scanning scheme maintains a balance between local detail and broader context.
Unlike LocalMamba, which relies on attention-based dynamic routing and suffers from unstable inference and high computational cost, LQMamba adopts a hierarchical architecture with fixed-size windows that change with network depth. As illustrated in Figure 1 (b), this structure enables the model to capture multi-scale perceptual cues, from fine-grained distortions to broader contextual patterns, while maintaining stable and efficient inference. It achieves a good balance between accuracy and efficiency, making it well-suited for IQA tasks that demand consistent perception across various distortion types and spatial scales.
Figure 3: t-SNE Visualization of Distortion-Specific Feature Separation: QMamba vs. Conventional Backbones
3.1.3 Analysis of Perceptual Capability in State Space Models
To investigate the feature selection mechanism of State Space Models (i.e.,Mamba) in visual quality assessment tasks, we employ t-SNE visualization for deep-layer feature analysis, as illustrated in Figure 3. Visualization results reveal that Mamba exhibits distinctive characteristic evolution patterns through dynamic state updating mechanisms, progressively enhancing distortion semantic perception across network hierarchies. Specifically, shallow layers preserve diverse features of original visual signals, while deeper layers adaptively filter redundant background information through gated state selection, intensifying distortion-type-specific feature focusing. This drives distortion-homogeneous samples to form cluster structures in feature space while amplifying inter-class discriminability. Such feature refinement strategy not only retains critical discriminative information for quality assessment but also significantly improves model sensitivity to quality degradation cues, thereby strengthening prediction robustness.